# DHG-Bench: A Comprehensive Benchmark on Deep Hypergraph Learning

### v1

@@ 4.3 Robustness Evaluation (RQ3)
To investigate the robustness of HNN algorithms, we focus on three key aspects: structural robustness, feature robustness, and supervision signal robustness. While our experiments primarily focus on the node classification task due to space constraints, the DHG-Bench library supports flexible extension to other tasks. We include 8 representative models from the three HNN categories.
4.3.1 Robustness analysis with respect to structure perturbations. To analyze structure-level robustness, we randomly remove or add node-hyperedge incidence connections in the original hypergraph structures of Cora and Actor, and evaluate the performance of HNN algorithms on the perturbed hypergraphs. The modification ratio is varied from 0 to 0.9 to simulate different levels of attack intensity. From Figure 3, we find that: (1) As the ratio of deletions or additions increases, model performance generally declines; however, the extent of degradation varies significantly across datasets. In particular, the accuracy drop caused by structural perturbations is more pronounced on the homophilic hypergraph Cora than on the heterophilic hypergraph Actor. (2) Spectral-based methods tend to be more susceptible to structural noise. For example, on the Cora dataset, as the deletion ratio increases from 0 to 0.9, spectral-based methods such as HGNN, SheafHyperGNN, and DPHGNN suffer an accuracy drop of over 30%, while methods from the other two categories consistently experience no more than a 10% decline.
4.3.2 Robustness analysis with respect to feature perturbations. To study feature-level robustness, we consider feature noise and sparsity. For feature noise, Gaussian noise is injected based on the dimensionality of node features. Sparsity is introduced by randomly masking a certain proportion of node features with zeros. As shown in Figure 4, we can observe that: (1) Although both increased feature noise and greater sparsity lead to performance degradation, sparsity has a more pronounced negative impact on accuracy than random feature noise. Specifically, on the Actor dataset, when the sparsity ratio increases from 0 to 0.9, all methods suffer an accuracy drop of about 10%. In contrast, under the same increase in noise level, 7 out of 8 methods experience accuracy losses of less than 5%. (2) The robustness of most algorithms to feature sparsity varies significantly across different datasets. In particular, most HNNs show better robustness on Actor than on Cora. At a feature mask ratio of 0.9, all algorithms on Cora suffer a performance drop exceeding 20%, whereas on Actor, more than half of the models show a decrease of less than 13%.
4.3.3 Robustness analysis with respect to supervision signals. On the basis of exploring feature and structural robustness, we further study supervision signal robustness by simulating realistic noise and sparsity scenarios. For label noise, a certain proportion of training samples are randomly assigned incorrect labels, while sparsity is introduced by varying the proportion of training nodes. Figure 5 indicates several key observations: (1) Model performance consistently declines as noise intensity increases or supervision signals become sparser, demonstrating a clear downward trend. Specifically, across different datasets, label noise leads to an approximately linear decline in accuracy as the noise ratio increases. In contrast, the effect of supervision sparsity is relatively modest at low sparsity levels but becomes markedly more severe at higher ratios, resulting in accelerated performance degradation. (2) Label noise has a more pronounced impact than feature-level or structural-level noise. Under the same noise level, label noise consistently causes greater performance degradation compared to feature or structure perturbations. For instance, when the label noise ratio reaches 15%, all HNN methods suffer an accuracy drop exceeding 15%. In contrast, most methods do not experience comparable degradation even when the feature or structural noise reaches 90%.
@@ 4.4 Fairness Evaluation (RQ4)
| Method | German | Bail | Credit | ||||||
| Acc | Acc | Acc | |||||||
| HJRL | 69.923.46 | 3.522.70 | 3.051.82 | OOM | OOM | OOM | OOM | OOM | OOM |
| ED-HNN | 70.163.15 | 4.063.05 | 4.072.75 | 94.260.77 | 8.080.64 | 1.530.26 | OOM | OOM | OOM |
We evaluate the fairness of various HNN algorithms on three datasets containing sensitive attributes. Predictive performance is measured by accuracy, while fairness is assessed using demographic parity () and equalized odds (). The experimental results in Table 4 show that: (1) While most existing HNN algorithms demonstrate stronger predictive performance, they tend to show more significant fairness concerns than traditional MLPs that do not rely on message passing. For instance, on the Credit dataset, MLP achieves of 3.43 and of 2.07, both lower than those of UniGNN, the fairest HNN model. This suggests that hypergraph message passing may exacerbate fairness issues in high-order representation learning. (2) Algorithmic fairness varies significantly across different datasets and evaluation metrics. No HNN method consistently demonstrates superior fairness across different settings. For instance, while HCHA achieves the best fairness performance on the Credit dataset across both metrics, its and rank as the second and third worst, respectively, on the German dataset. Moreover, the results reveal that and capture different aspects of fairness, and better performance in one metric does not necessarily indicate better performance in the other. For example, on the Bail dataset, although HNHN attains the lowest among all HNN methods, its ranks as the third highest among the 16 models. These observations highlight the necessity of developing fairness-aware HNN algorithms that can achieve robust fairness across diverse datasets and evaluation metrics.
5 Conclusion and Future Directions
@@
This paper presents DHG-Bench, the first comprehensive benchmark for deep hypergraph learning, which integrates and compares 16 representative HNNs across 20 diverse hypergraph datasets encompassing various domains, sizes, and structural properties. We conduct extensive experiments to evaluate the effectiveness, efficiency, robustness, and fairness of DHGL algorithms on node-level, edge-level, and graph-level tasks. Our analysis reveals the strengths and weaknesses of different HNN models under a wide range of settings, offering valuable insights into their practical applicability and design trade-offs. Furthermore, we develop and release a package DHG-Bench (https://github.com/Coco-Hut/DHG-Bench) that includes all experimental protocols, baseline algorithms, datasets, and reproducibility scripts to facilitate future research.
Drawing upon our empirical analyses, we point out some promising future directions for DHGL community.
Designing a unified algorithm for diverse tasks and datasets. Our experiments reveal that existing HNN architectures show substantial performance disparities across datasets and tasks, which hinders their transferability to diverse scenarios. Future research should explore novel architectural designs and training strategies that promote robust generalization across heterogeneous datasets and tasks, ultimately moving toward a versatile “one-for-all” solution in DHGL.
Improving the Efficiency and Scalability of HNN methods. Our results show that many HNN algorithms either fail to scale to large datasets or suffer significant performance degradation when applied to them. Given that real-world hypergraphs in industrial applications often contain millions of nodes and hyperedges, future work should focus on designing HNN algorithms that deliver both high efficiency and strong effectiveness at scale. Progress in this direction will be essential for enabling the practical adoption of HNNs in industrial settings.
Enhancing the robustness of HNN algorithms. Our experimental results show that current HNN methods are susceptible to data perturbations across multiple aspects, including structural modifications, feature noise, and corrupted supervisory signals. Future research should focus on designing robust HNN algorithms that are resilient to such perturbations and maintain reliable performance in noisy industrial environments.
Developing fairness-aware HNN algorithms. Empirical evidence suggests that current message passing-based HNN algorithms often suffer from fairness issues, leading to biased predictions that disadvantage certain demographic groups or individuals. Future research should explore how to develop fairness-aware HNN methods that mitigate such biases and discrimination, enabling their safe adoption in high-stakes real-world applications such as crime prediction (Suresh & Guttag, 2019), job hunting (Mehrabi et al., 2021), and credit evaluation (Yeh & Lien, 2009).
@@ Appendix A Datasets and Algorithms
A.1 Detailed description of datasets
We adopt 20 publicly available benchmark datasets, ranging from node-level to graph-level, to comprehensively evaluate HNN algorithms in terms of effectiveness, efficiency, robustness, and fairness. Detailed descriptions of these datasets are provided below.
Node/Edge-level Datasets
Cora/Pubmed/Cora-CA/DBLP-CA (Yadati et al., 2019): Cora and Pubmed are co-citation networks where nodes represent papers and hyperedges connect papers cited together. Cora-CA and DBLP-CA are co-authorship hypergraphs, with nodes as papers and hyperedges linking all papers co-authored by the same author. Node features are Bag-of-Words representations of the documents, and labels indicate paper categories.
Warmart (Chien et al., 2022): The Walmart dataset models a hypergraph where nodes represent products and hyperedges capture sets of products purchased together. Node labels indicate product categories. Following (Chien et al., 2022), each node feature is a 100-dimensional vector obtained by adding Gaussian noise with to one-hot encodings of the labels.
Trivago (Kim et al., 2023): Trivago is a hotel-web search hypergraph where each node indicates each hotel and each hyperedge corresponds to a user. If a user (hyperedge) has visited the website of a particular hotel (node), the corresponding node is added to the respective user hyperedge. Furthermore, each hotel’s class is labeled based on the country in which it is located.
Actor (Li et al., 2025): The actor co-occurrence network is derived from a heterogeneous movie-actor-director-writer network 222https://www.aminer.org/lab-datasets/soinf/, capturing intricate collaborations within films. Nodes represent individuals involved in film production (actors, directors, writers), and hyperedges denote their joint participation in a single film. Node attributes are extracted from Wikipedia keywords, and labels indicate each individual’s specific role.
Amazon-ratings (Ratings) (Li et al., 2025): This dataset, sourced from the Amazon co-purchasing network in the SNAP repository (Leskovec, 2014), includes products like books, music CDs, DVDs, and VHS tapes. Nodes represent individual products, and hyperedges link those frequently co-purchased together. The task is to predict each product’s average user rating, classified into ten levels. Node features are extracted using the Bag of Words method applied to product descriptions (Juluru et al., 2021).
Twitch-gamers (Gamers) (Li et al., 2025): The Twitch-gamers dataset is a connected undirected hypergraph representing user interactions on the Twitch streaming platform. Nodes denote user accounts, and hyperedges are formed based on mutual follows within specific timeframes. Each node is associated with features such as view counts, timestamps, language preferences, activity duration, and inactivity status. The goal is to predict whether a channel hosts explicit content (binary classification).
Pokec (Li et al., 2025): The Pokec dataset is derived from Slovakia’s largest online social networking platform and is used to model social relationships and attributes. Nodes represent individual users, and hyperedges correspond to each user’s full set of friends. Node labels indicate user-reported gender, while node features are extracted from profile information, including age, hobbies, interests, education level, region, etc.
Yelp (Chien et al., 2022): The Yelp dataset is a hypergraph where nodes represent restaurants and hyperedges link those visited by the same user. Node labels denote average star ratings (1.0–5.0 in 0.5 steps). Features include geographic coordinates, one-hot encodings of city/state, and Bag-of-Words vectors from the top-1000 restaurant name tokens.
German Credit (German) (Wei et al., 2022): The nodes in the dataset represent clients in German Bank, and hyperedges are constructed by linking individuals with the most similar credit accounts to each person in the dataset. The task is to classify credit risk levels as high or low based on the sensitive attribute ”gender” (Male/Female).
Credit Defaulter (Credit) (Wei et al., 2022): The nodes in the dataset represent credit card users, and hyperedges are formed based on the similarity of users’ spending and payment patterns. The task is to classify the default status with the sensitive attribute ”age” (¡25/¿25).
Graph-level Datasets
RHG-10/RHG-3 (Feng et al., 2024): RHG-10 dataset encompasses ten distinct synthetic factor hypergraph structures (i.e., Hyper Flower, Hyper Pyramid, Hyper Checked Table, Hyper Wheel, Hyper Lattice, Hyper Windmill, Hyper Firm Pyramid, Hyper RChecked Table, Hyper Cycle, and Hyper Fern). To evaluate the algorithm’s ability to recognize significant high-order structures, RHG-3 dataset is constructed by randomly generating hypergraphs for three distinctively various hypergraph structures: hyper-pyramid, hyper-check-table, and hyper-wheel.
IMDB-Dir-Form/IMDB-Dir-Genre (Feng et al., 2024): These two datasets contain hypergraphs constructed by the co-director relationship from the original IMDB dataset. The director of each movie is a hypergraph. ”Form” included in the dataset’s name indicates that the movie category is identified by its form, like animation, documentary, and drama. ”Genre” denotes the movie is classified by its genres, like adventure, crime, and family.
Stream-Player (Feng et al., 2024): The Steam-Player dataset is a player dataset where each player is a hypergraph. The vertex is the games played by the player, and the hyperedge is constructed by linking the games with shared tags. The target of the dataset is to identify each user’s preference: single-player-game or multi-player-game.
| Datasets | Nodes | Edges | Features | Classes | Type | Description |
| Cora | 2,708 | 1,579 | 1,433 | 7 | Homophily | Co-citation |
| PubMed | 19,717 | 7,963 | 500 | 3 | Homophily | Co-citation |
| Cora-CA | 2,708 | 1,072 | 1,433 | 7 | Homophily | Co-author |
| DBLP-CA | 41,302 | 22,363 | 1,425 | 6 | Homophily | Co-author |
| Warmart | 88,860 | 69,906 | 100 | 11 | Homophily | Co-purchasing |
| Trivago | 172,738 | 233,202 | 300 | 160 | Homophily | Co-click |
| Actor | 16,255 | 10,164 | 50 | 3 | Heterophily | Co-occurence |
| Ratings | 22,299 | 2,090 | 111 | 5 | Heterophily | Co-purchasing |
| Gamers | 16,812 | 2,627 | 7 | 2 | Heterophily | Co-create |
| Pokec | 14,998 | 2,406 | 65 | 2 | Heterophily | Co-friendship |
| Yelp | 50,758 | 679,302 | 1,862 | 9 | Heterophily | Co-visit |
| Datasets | German Credit | Bail | Credit Defaulter |
| Nodes | 1000 | 18876 | 30000 |
| Edges | 1000 | 18876 | 30000 |
| Features | 27 | 18 | 13 |
| Sens Attribute | Gender | Race | Age |
| Labels | Credit status | Bail decision | Future default |
| Datasets | Hypergraphs | Avg. | Avg. | Avg. | Classes |
@@ A.1 Detailed description of datasets
A.2 Detailed description of algorithms

### v2

@@ 4.3 Robustness Evaluation (RQ3)
In this section, we assess HNN robustness by simulating structural, feature, and supervision perturbations, as detailed in Appendix B.4. While our experiments primarily focus on the node classification task due to space limits, DHG-Bench supports flexible extension to other tasks. We evaluate 10 representative models on four datasets (Cora, Pubmed, Actor, and Pokec). The results on Pubmed and Pokec (Figures A2, A3, and A4) are provided in Appendix C.2.
4.3.1 Robustness Analysis with respect to Structure Perturbations
Results (Figure 2 and Figure A2). Most HNN algorithms exhibit strong robustness against random structural noise, experiencing only marginal performance drops or even remaining nearly unaffected under high perturbation rates. For example, when 90% of hyperlinks are randomly removed from Cora, 7 out of 10 methods degrade by less than 7%. Similarly, when 90% of random hyperlinks are injected into Actor, only 2 models show a noticeable decline in performance. Spectral-based approaches are generally more vulnerable to structural perturbations. On Pubmed, for instance, increasing the ratio of noisy hyperlinks results in a pronounced performance decline across four spectral-based methods (HGNN, PhenomNN, DPHGNN, and TF-HNN), whereas most other methods remain stable. This may be because spectral methods rely on the hypergraph’s global eigenstructure, which is highly sensitive to topological noise. The robustness of HNN algorithms varies with both the type of structural perturbation (deletion vs. addition) and the choice of dataset. For example, on the Actor dataset, SheafHyperGNN suffers substantial performance degradation under hyperlink deletion but demonstrates strong robustness under hyperlink addition. In another case, PhenomNN exhibits strong robustness on Cora in the addition scenario while showing the opposite trend on Actor.
4.3.2 Robustness Analysis with respect to Feature Perturbations
Results (Figure 3 and Figure A3). Feature perturbations under equal noise or sparsity levels result in greater performance degradation than structural ones, indicating a more critical role of node features in model prediction. With increasing noise intensity, model accuracy decreases sharply at the beginning and then stabilizes, as highly corrupted features approximate randomness and lose predictive utility. As the feature masking rate increases, model performance degrades progressively faster, with a slow decline at low ratios and a sharp drop under high sparsity. Compared to feature sparsity, feature noise poses a greater challenge for HNN algorithms, with equivalent levels of noise typically resulting in lower predictive accuracy across different datasets.
4.3.3 Robustness Analysis with respect to Supervision Perturbations
Results (Figure 4 and Figure A4). As noise intensity increases or supervision becomes sparser, all models show a clear downward trend in performance, with label noise exerting a more pronounced impact. Increasing label noise generally causes a rapid yet steady decline in performance, which appears approximately linear in most cases. The impact of supervision sparsity is modest at lower levels but intensifies at higher ratios, resulting in an accelerating decline in model performance. This trend highlights the challenges faced by current HNNs in low-label scenarios. Label noise and sparsity tend to degrade performance more substantially on homophilic datasets than on heterophilic ones, reflecting the reliance of model predictions on data homophily.
@@ 4.4 Fairness Evaluation (RQ4)
In this section, we analyze algorithmic fairness and report full quantitative results in terms of accuracy (Acc), , and in Table A8 of Appendix C.3. To better illustrate the strengths and limitations of each algorithm, we present Figure 5, which shows their average rankings across the three metrics on datasets where they can run, considering only HNNs executable on at least two datasets.
Results (Figure 5 and Table A8). While HNN algorithms achieve higher predictive performance, they generally suffer from more severe fairness issues compared to MLP, which is free from message passing. Figure 5 shows that MLP ranks best on the two fairness metrics but worst on accuracy. For example, on the Credit dataset, MLP achieves lower and values than HCHA, the fairest among the evaluated HNN models, as shown in Table A8. The fairness performance of HNN algorithms varies considerably across datasets, with no method achieving consistently superior performance on all benchmarks. For instance, Table A8 illustrates that while HCHA achieves the best fairness performance on the Credit dataset across both metrics, its and rank as the second- and third-worst, respectively, on the German dataset. Moreover, Figure 5 shows that most algorithms exhibit substantial variance in their rankings, further highlighting the instability of fairness across datasets. HNN algorithms show inconsistent behavior across fairness metrics, and strong performance on one does not guarantee superiority on another. For example, on the Bail dataset, although HNHN achieves the lowest among all HNN methods, its ranks as the third worst among the 17 HNN models.
5 Conclusion
@@
This paper introduces DHG-Bench, the first comprehensive benchmark for deep hypergraph learning, which integrates and compares 17 representative HNNs across 22 hypergraph datasets encompassing various domains, sizes, and structural properties, under consistent experimental settings. We comprehensively evaluate the effectiveness, efficiency, robustness, and fairness of HNN algorithms, and our analysis reveals the strengths and weaknesses of different HNNs in a wide range of scenarios, offering valuable insights into their practical applicability and design trade-offs. Furthermore, we develop and release a package, DHG-Bench, that includes all experimental protocols, baseline algorithms, datasets, and reproducibility scripts to facilitate future research. To foster further advances in deep hypergraph learning, we outline several promising future directions for HNNs in Appendix D.2.
Appendix
@@ Appendix A Datasets and Algorithms
A.1 Benchmark Datasets
| Dataset | # Nodes | # Edges | # Features | Avg. | # Classes | |
| Cora | 2,708 | 1,579 | 1,433 | 3.03 | 0.75 | 7 |
| Pubmed | 19,717 | 7,963 | 500 | 4.35 | 0.78 | 3 |
| Cora-CA | 2,708 | 1,072 | 1,433 | 4.28 | 0.78 | 7 |
| DBLP-CA | 41,302 | 22,363 | 1,425 | 4.45 | 0.87 | 6 |
| NTU2012 | 2,012 | 2,012 | 100 | 5.00 | 0.79 | 67 |
| ModelNet40 | 12,311 | 12,311 | 100 | 5.00 | 0.87 | 40 |
| Walmart | 88,860 | 69,906 | 100 | 6.59 | 0.60 | 11 |
| Trivago | 172,738 | 233,202 | 300 | 3.12 | 0.98 | 160 |
| Actor | 16,255 | 10,164 | 50 | 5.25 | 0.46 | 3 |
| Ratings | 22,299 | 2,090 | 111 | 3.10 | 0.37 | 5 |
| Gamers | 16,812 | 2,627 | 7 | 6.23 | 0.49 | 2 |
| Pokec | 14,998 | 2,406 | 65 | 2.29 | 0.45 | 2 |
| Yelp | 50,758 | 679,302 | 1,862 | 6.66 | 0.29 | 9 |
| Dataset | # Nodes | # Edges | # Features | Sens | Label |
| German | 1,000 | 1,000 | 27 | Gender | Credit status |
| Bail | 18,876 | 18,876 | 18 | Race | Bail decision |
| Credit | 30,000 | 30,000 | 13 | Age | Future default |
| Dataset | # Hypergraphs | Avg. | Avg. | Avg. | # Classes |
@@ A.1 Benchmark Datasets
We adopt 22 publicly available benchmark datasets to comprehensively evaluate HNN algorithms. The statistics of node-level datasets, fairness-sensitive datasets, and graph-level datasets are reported in Tables A1, A2, and A3, respectively. Detailed descriptions of these datasets are provided below.
Cora/Pubmed/Cora-CA/DBLP-CA (Yadati et al., 2019): Cora and Pubmed are co-citation networks where nodes represent papers and hyperedges connect papers cited together. Cora-CA and DBLP-CA are co-authorship hypergraphs, with nodes as papers and hyperedges linking all papers co-authored by the same author. Node features are Bag-of-Words (BoW) (Zhang et al., 2010) representations of the documents, and labels indicate paper categories.
NTU2012/ModelNet40 (Feng et al., 2019): The ModelNet40 and the NTU2012 are two computer vision and graphics datasets. ModelNet40 contains 12,311 3D objects from 40 popular categories, while NTU2012 consists of 2,012 3D shapes from 67 categories. For each object, features are extracted using both the Group-View Convolutional Neural Network (GVCNN)(Feng et al., 2018) and the Multi-View Convolutional Neural Network (MVCNN)(Su et al., 2015). Following (Feng et al., 2019), we construct hyperedges by aggregating the nearest neighbors of each node based on Euclidean distance.
Walmart (Chien et al., 2022): The Walmart dataset models a hypergraph where nodes represent products and hyperedges capture sets of products purchased together. Node labels indicate product categories. Following (Chien et al., 2022), each node feature is a 100-dimensional vector obtained by adding Gaussian noise with to one-hot encodings of the labels.
Trivago (Kim et al., 2023): Trivago is a hotel-web search hypergraph where each node indicates a hotel, and each hyperedge corresponds to a user. If a user (hyperedge) has visited the website of a particular hotel (node), the corresponding node is added to the respective user hyperedge. Furthermore, each hotel’s class is labeled based on the country in which it is located.
Actor (Li et al., 2025c): The actor co-occurrence network is derived from a heterogeneous movie-actor-director-writer network 222https://www.aminer.org/lab-datasets/soinf/, capturing intricate collaborations within films. Nodes represent individuals involved in film production (actors, directors, and writers), and hyperedges denote their joint participation in a single film. Node attributes are extracted from Wikipedia keywords, and labels indicate each individual’s specific role.
Amazon-ratings (Ratings) (Li et al., 2025c): This dataset, sourced from the Amazon co-purchasing network in the SNAP repository (Leskovec, 2014), includes products like books, music CDs, DVDs, and VHS tapes. Nodes represent individual products, and hyperedges link those frequently purchased together. The task is to predict each product’s average user rating, classified into ten levels. Node features are extracted using the BoW technique applied to product descriptions (Juluru et al., 2021).
Twitch-gamers (Gamers) (Li et al., 2025c): The Twitch-gamers dataset is a connected undirected hypergraph representing user interactions on the Twitch streaming platform. Nodes denote user accounts, and hyperedges are formed based on mutual follows within specific timeframes. Each node is associated with features such as view counts, timestamps, language preferences, activity duration, and inactivity status. The goal is to predict whether a channel hosts explicit content (binary classification).
Pokec (Li et al., 2025c): The Pokec dataset is derived from Slovakia’s largest online social networking platform and is used to model social relationships and attributes. Nodes represent individual users, and hyperedges correspond to each user’s full set of friends. Node labels indicate user-reported gender, while node features are extracted from profile information, including age, hobbies, interests, education level, region, etc.
Yelp (Chien et al., 2022): The Yelp dataset is a hypergraph where nodes represent restaurants and hyperedges link those visited by the same user. Node labels denote average star ratings (1.0–5.0 in 0.5 steps). Features include geographic coordinates, one-hot encodings of city/state, and BoW vectors from the top-1000 restaurant name tokens.
German (Wei et al., 2022): The nodes in the dataset represent clients in the German Bank, and hyperedges are constructed by linking individuals with the most similar credit accounts to each person in the dataset. The task is to classify credit risk levels as high or low based on the sensitive attribute ”gender” (Male/Female).
Credit (Wei et al., 2022): The nodes in the dataset represent credit card users, and hyperedges are formed based on the similarity of users’ spending and payment patterns. The task is to classify the default status with the sensitive attribute ”age” (25 / 25).
RHG-10/RHG-3 (Feng et al., 2024): RHG-10 dataset encompasses ten distinct synthetic factor hypergraph structures (i.e., Hyper Flower, Hyper Pyramid, Hyper Checked Table, Hyper Wheel, Hyper Lattice, Hyper Windmill, Hyper Firm Pyramid, Hyper RChecked Table, Hyper Cycle, and Hyper Fern). To evaluate the algorithm’s ability to recognize significant high-order structures, the RHG-3 dataset is constructed by randomly generating hypergraphs for three distinctively various hypergraph structures: Hyper Pyramid, Hyper Checked Table, and Hyper Wheel.
IMDB-Dir-Form/IMDB-Dir-Genre (Feng et al., 2024): These two datasets contain hypergraphs constructed by the co-director relationship from the original IMDB dataset. The director of each movie is a hypergraph. ”Form” included in the dataset’s name indicates that the movie category is identified by its form, like animation, documentary, and drama. ”Genre” denotes that the movie is classified by its genres, like adventure, crime, and family.
Steam-Player (Feng et al., 2024): The Steam-Player dataset is a player dataset where each player is a hypergraph. The vertex is the games played by the player, and the hyperedge is constructed by linking the games with shared tags. The target of the dataset is to identify each user’s preference: single-player game or multiplayer game.
A.2 Benchmark Algorithms